OpenStudy (anonymous):

When solving negative one over eight (x + 35) = −7, what is the correct sequence of operations? Multiply each side by negative one over eight , add 35 to each side Multiply each side by negative one over eight , subtract 35 from each side Multiply each side by −8, subtract 35 from each side Multiply each side by −8, add 35 to each side

OpenStudy (anonymous):

So do you care if I teach you order of operations? It is very important to know.

OpenStudy (anonymous):

go ahead

OpenStudy (anonymous):

c

OpenStudy (anonymous):

P.E.M.D.A.S., ever heard of it? P - Parenthesis E - Exponents M - Multiply D - Divide A - Add S - Subtract This is the order of operations. A easy way to remember it is by saying the phrase "Please excuse my dear Aunt Sally."

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

So can you tell me what you think the answer is? Give me a reason as well so I know your not just guessing.

OpenStudy (anonymous):

i think it is C because when i started working it out the other answer have you doing 1/8 first though you dont do that yet

OpenStudy (anonymous):

If you are still struggling I can understand, this is a tricky question. The work is shown below: ( x + 35 ) = -7 35 - ( x + 35 ) = -7 - 35 x = -42 I would say all the answers are wrong, however they are looking for C.

OpenStudy (anonymous):

okay thank you so much!